### v1

@@ 2.3 Scientific Workflows
Modern scientific collaborations have opened up the opportunity to solve complex problems that demand multidisciplinary expertise and large-scale computational experimentation. These experiments typically involve a sequence of processing steps that need to be executed on selected computing platforms. However, executing these experiments can be difficult due to the complexity and variety of applications, the diversity of analysis goals, the heterogeneity of computing platforms, and the volume and distribution of data. A common strategy to make these experiments more manageable is to model them as workflows and to use an SWS to organize their execution [59]. Lin et al. [24] defined scientific workflow as a computerized facilitation or automation of a scientific process, in whole or part, which usually streamlines a collection of scientific tasks with data channels and dataflow constructs to automate data computation and analysis to enable and accelerate scientific discovery.
@@
We show a script-based (DSL) workflow using Snakemake SWS that explores the Iris dataset from Kaggle in Figure 3. This workflow downloads the dataset, performs basic exploratory data analysis (EDA), and generates a summary report.
@@ 2.4 Topic Modeling
Topic modeling is a statistical modeling technique used to uncover the underlying structure and main themes within a large collection of documents, making it easier to organize, search, and understand the content [63]. Its applications span various text analysis tasks [64], including document clustering [65], information retrieval [66], recommendation systems [67], language translation [68], and content summarization [69]. Researchers have extensively explored topic modeling, publishing numerous articles in software engineering, political science, medicine, and linguistics. Extensive utilization of topic modeling is observed in the software engineering research community [70], particularly in domains like mining software repositories [71, 72, 73, 74, 75, 76, 77, 78], source code analysis [79, 80, 81, 82, 83, 78], spam detection [84], and recommendation systems [85, 86, 87, 88, 89], among others. While numerous papers contribute to this field, noting all of them is beyond our scope. Hence, we highlight a selection of significant papers. Several methods, such as BERTopic [90], Latent Dirichlet Allocation (LDA) [91], Non-Negative Matrix Factorization (NMF) [92], Latent Semantic Analysis (LSA) [93], Correlated Topic Model (CTM) [94], Dynamic Topic Models [95], Word Embedding-Based Models [95], and Biterm Topic Model (BTM) [96], contribute to the advancements of topic modeling.
@@ 2.5 Topic Analysis of Technical Q&As
Topic modeling has significant applications in technical Q&A data, facilitating the analysis and categorization of large volumes of questions and answers. By identifying underlying themes and topics within the content, topic modeling helps to organize and index data, making it easier for users to find relevant information. For instance, in forums such as Stack Overflow, topic modeling can cluster similar questions, aiding in quicker retrieval of solutions and reducing redundancy [102]. Additionally, it can track trending topics, monitor user interests, and identify common issues or knowledge gaps, proving valuable for both users seeking information and developers aiming to improve their products or services [103]. Topic models are extensively used in many studies to understand the topics of general SO posts and track topic trends [104, 105, 106, 107]. Prior work also leverages topic modeling in specific application development domains, such as mobile application development [16], machine learning application development [108], concurrency development [109], security-related development [20], DevOps development [110], chatbot development [17], classification of database technology [111], and understanding nonfunctional requirements [112]. To the best of our knowledge, topic analysis of SWSs-related posts on SO is still pending, but SWS is a promising field of software engineering. Therefore, we recognize the need to identify the topics of SWSs, and we utilize BERTopic [90] because of its contextual understanding capacity and high accuracy.
@@ 2.6 Topic Analysis of GitHub Data
Topic modeling with GitHub data involves analyzing text from repositories, issues, pull requests, and comments to identify prevalent themes or topics. This technique can uncover trends, common issues, or popular features within software development communities. Researchers apply algorithms like LDA, BERTopic, and others to categorize the vast amount of unstructured text data on GitHub, helping developers and organizations better understand user needs, project focus, and collaboration patterns. Previous studies have utilized topic modeling to analyze GitHub data, revealing prominent challenges developers face, user requirements, and emerging trends in software projects. This insight assists project managers in task prioritization and resource allocation while promoting streamlined collaboration and well-informed decision-making. By delving into the dynamics of software development communities, researchers contribute to improving software engineering methodologies. For example, Scoccia et al. [22] used LDA topic modeling to identify challenges faced by web developers, correlating real-time GitHub issues with queries on SO to explore common areas in web app development. Similarly, Li et al. [21] applied LDA topic modeling to understand challenges in quantum software engineering by analyzing SO posts and GitHub issue reports. Several other studies [113, 114, 115, 75, 116, 117] have utilized topic modeling with GitHub datasets. Given that GitHub issues and pull requests are often brief and context-specific, BERTopic’s embedding-based approach is particularly well-suited for analyzing this data. Therefore, we choose to utilize BERTopic [90] for our analysis.
@@ 3 Study Design
The main goal of our study is to examine what SWSs developers are asking about. To achieve this goal, we analyze discussions on SO posts, and issues and pull requests on GitHub. These platforms provide rich datasets, frequently used in research across various domains, including quantum software engineering [21], cryptography APIs [118], concurrency [109], and deep learning [119]. SO offers structured data with questions, answers, and metadata (e.g., accepted answers, scores) but lacks fine-grained topic information related to SWSs. Hence, we first need to identify posts from SO related to SWSs, group them according to their dominant topic, and then conduct our analysis. For GitHub, we focus on projects related to SWSs identified from our SO data, ensuring consistency across both platforms in the projects we analyze. Figure 4 shows the overall study design of our research. Below, we describe the study design approaches of our work.
@@ 3.1 Stack Overflow Data Collection and Prepossessing
Stack Overflow is a widely used technical Q&A forum among developers. Therefore, we chose it as one of our data sources. Below, we outline the steps involved in collecting and preprocessing SO data.
Step 1: Gather Scientific Workflow Systems List. Researchers typically use tag-based search functionality on SO to retrieve data, as authors tag their posts with tags (e.g., javascript, python, arrays) to enhance visibility and increase the possibility of receiving answers [120]. However, the field of SWSs spans multiple disciplines (e.g., bioinformatics [121], geospatial analysis [122], software engineering [123], machine learning [124], grid computing [125]), leading to diverse terminologies used by different users. This terminological inconsistency presents a challenge when attempting to conduct searches related to SWSs on SO. Additionally, SWSs have different names (e.g., Galaxy [27], NextFlow [34], SnakeMake [33], Dask [126]) depending on the specific domain or community, further complicating the process of locating these systems through specific tags on SO. Thus, we plan to gather the SWSs list and use it to identify the posts related to them. We compiled a list of SWSs by referencing multiple sources [127, 128, 129, 130]. Upon merging these lists, we identified overlaps, necessitating the removal of duplicates to ensure the integrity and completeness of our compilation. We also added the web link and GitHub link of each SWS wherever possible. Finally, we listed 352 SWS, which can be accessed using the replication package [51].
@@
We took precautions against potentially noisy data to ensure the accuracy of our dataset and remove any irrelevant information. Despite the possibility of encountering posts with the term celery in their title or body unrelated to the celery workflow system, we conducted a meticulous filtering process. By initially narrowing down our dataset using celery tags, we identified 9,499 posts specifically discussing the celery workflow system. This validation was reinforced by cross-referencing the tags’ descriptions provided by SO. Subsequently, we were left with 139 posts, each potentially containing content unrelated to the celery SWS. These posts varied, with some mentioning the term celery in contexts unrelated to our focus, such as referring to the vegetable, a project name, a part of a filename, or a list ID. To meticulously ensure the precision of our dataset, each of these remaining 139 posts underwent manual scrutiny. Initially, one of the authors, having over four years of experience with SWSs, meticulously reviewed each post, categorizing them as either related or unrelated to the celery workflow system. Following this initial review, two authors with over eight years of experience further scrutinized the posts. Only upon reaching a consensus did we include any posts in our final dataset. This rigorous process enabled us to identify and exclude six posts unrelated to the celery workflow system (provided in the replication package). Consequently, we finalized our dataset with 9622 posts exclusively pertinent to the celery workflow system.
@@
In some cases, we encountered posts where all content was directly pertinent to SWSs. For instance, in the case of SnakeMake, our extraction process yielded a total of 1,696 posts, meticulously considering tags, Title, and Body. Remarkably, 1,695 of these posts were tagged with snakemake SWS, solidifying their relevance to the SnakeMake SWS. The one remaining post not tagged with SnakeMake also discussed the SnakeMake SWS, further affirming its alignment with our focus. Similar outcomes were observed for KNIME. A thorough investigation led us to identify a grand total of 308 posts related to KNIME, with 305 of them bearing the knime tags. The remaining three, lacking the knime tags, were inherently related to KNIME, underscoring their inclusion in our dataset.
@@
Utilizing similar approaches, we explored the Galaxy SWS related posts. Our search encompassed Tags, Title, and Body, yielding a substantial total of 5652 posts. However, upon closer examination, we noted many posts tagged with android that were unrelated to Galaxy SWS. Consequently, we opted to discard these posts, amounting to the exclusion of 4378 entries. Subsequently, from the remaining 1274 posts, we further refined our dataset by filtering out posts tagged with samsung, mobile, or containing these terms in their title and body. This process led to the removal of 510 additional posts. With 764 posts remaining, we then focused specifically on those related to Galaxy SWS by filtering with the term ansible, associated with Galaxy. This step resulted in the identification of 165 posts relevant to Galaxy. Continuing our scrutiny, we observed 26 posts concerning PHP, 35 posts discussing Java, and 77 posts referencing Galaxy S4-S7. These posts were deemed irrelevant to Galaxy SWS and were consequently excluded. From the remaining 461 posts, one author meticulously assessed each post, determining their relevance to Galaxy SWS. Subsequently, the other two authors verified these assessments. Only posts that were unanimously agreed upon as relevant to Galaxy SWS were retained. Ultimately, this rigorous vetting process yielded ten posts directly related to Galaxy SWS. Thus, our finalized dataset comprises a total of 175 posts aligned with Galaxy SWS.
@@
1 Airflow 10602 10589 33163
2 SnakeMake 1696 1696 6187
3 Celery 9628 9622 13302
4 NextFlow 419 418 5397
5 KNIME 308 308 153
6 Galaxy 5648 175 39887
7 Oozie 2066 2052 674
8 Drake 846 846 20791
9 Luigi 362 348 3269
10 Dask 4681 4662 28611
11 NiFi 5187 5187 8440
For airflow, we obtained 10602 posts and we found 10448 posts with airflow tags and 31 posts having apache in title and body. Then, from the remaining 123 posts we manually checked each posts and found 13 posts unrelated with Airflow SWS. They are mainly talking about heat transfer problems, airflow temperature, respiratory airflow, function name etc. Thus, finally we have 10589 posts related with Airflow SWS.
As there are hundreds of SWS, describing each SWS is beyond our scope, though our approach is identical to that described above. In Table 1, we show the SWSs, total posts obtained after the query, and total filtered posts. For simplicity, we display the SWSs where we found more than 100 posts for a particular SWS. Our further analysis is also based on the dataset shown in Table 1. These systems are publicly available state-of-the-art systems designed for a broad spectrum of scientific analysis tasks. We gather a total of 41,443 posts. After filtration, we identified 5,540 posts unrelated to the scientific workflow system, and the remaining 35,903 posts are about SWS.
@@ 11 NiFi 5187 5187 8440
Step 3: Preprocessing Scientific Workflow Systems’ posts. We first filtered out the irrelevant information before applying topic modeling techniques. While the post Title concisely summarizes the question, the Body provides crucial context and details that help us accurately identify the topic being discussed. Therefore, in this analysis, we focused on both the post Title and Body. However, the Body content might introduce noise to our analysis, so we removed quotes, HTML tags, links, and any code snippets using regular expressions from the Body of the posts. We also remove stopwords, commonly used words in the English language, such as ’how’, ’can’, and ’at’ that do not significantly affect the meaning of a sentence and can introduce bias. We used the NLTK stopwords list [132] for this operation. Then, we performed lemmatization, which involves reducing a word to its canonical form called a lemma, taking into account the word’s linguistic context (for example, the word ’scale’ is the lemma of the word ’scaling’). All the data preprocessing steps and processed data are provided in the replication package [51].
3.2 GitHub Data Collection and Prepossessing

### v2

@@ 2.3 Scientific Workflows
Modern scientific collaborations have opened up the opportunity to solve complex problems that demand multidisciplinary expertise and large-scale computational experimentation. These experiments typically involve a sequence of processing steps that need to be executed on selected computing platforms. However, executing these experiments can be difficult due to the complexity and variety of applications, the diversity of analysis goals, the heterogeneity of computing platforms, and the volume and distribution of data. A common strategy to make these experiments more manageable is to model them as workflows and to use an SWS to organize their execution [61]. A workflow defines the sequence of activities, their dependencies, and the data flow between them, enabling researchers to automate, manage, and reproduce complex scientific computations [62]. It is akin to a cooking recipe, where ingredients represent datasets, and instructions describe the sequence of operations to produce a result (e.g., a plot or new dataset). The order of operations is important, as each step often uses the output of the previous one, similar to sifting flour before mixing it with eggs when baking a cake. This sequence of steps is referred to as a pipeline. Multiple pipelines can combine to form a workflow, much like a menu composed of various recipes. Workflows allow the consistent application of procedures to different datasets by changing the inputs, ensuring adaptability and reproducibility.
@@
We showed a script-based (DSL) workflow using Snakemake SWS that explores the Iris dataset from Kaggle in Figure 3. This workflow downloads the dataset, performs basic exploratory data analysis (EDA), and generates a summary report.
@@ 2.4 Topic Modeling
Topic modeling is a statistical modeling technique used to uncover the underlying structure and main themes within a large collection of documents, making it easier to organize, search, and understand the content [66]. Its applications span various text analysis tasks [67], including document clustering [68], information retrieval [69], recommendation systems [70], language translation [71], and content summarization [72]. Researchers have extensively explored topic modeling, publishing numerous articles in software engineering, political science, medicine, and linguistics. Extensive utilization of topic modeling is observed in the software engineering research community [73], particularly in domains like mining software repositories [74, 75, 76, 77, 78, 79, 80, 81], source code analysis [82, 83, 84, 85, 86, 81], spam detection [87], and recommendation systems [88, 89, 90, 91, 92], among others. While numerous papers contribute to this field, noting all of them is beyond our scope. Hence, we highlighted a selection of significant papers. Several methods, such as BERTopic [93], Latent Dirichlet Allocation (LDA) [94], Non-Negative Matrix Factorization (NMF) [95], Latent Semantic Analysis (LSA) [96], Correlated Topic Model (CTM) [97], Dynamic Topic Models [98], Word Embedding-Based Models [98], and Biterm Topic Model (BTM) [99], contribute to the advancements of topic modeling.
@@ 2.5 Topic Analysis of Technical Q&As
Topic modeling has significant applications in technical Q&A data, facilitating the analysis and categorization of large volumes of questions and answers. By identifying underlying themes and topics within the content, topic modeling helps to organize and index data, making it easier for users to find relevant information. For instance, in forums such as Stack Overflow, topic modeling can cluster similar questions, aiding in quicker retrieval of solutions and reducing redundancy [105]. Additionally, it can track trending topics, monitor user interests, and identify common issues or knowledge gaps, proving valuable for both users seeking information and developers aiming to improve their products or services [106]. Topic models are extensively used in many studies to understand the topics of general SO posts and track topic trends [107, 108, 109, 110]. Prior work also leverages topic modeling in specific application development domains, such as mobile application development [21], machine learning application development [111], concurrency development [112], security-related development [25], DevOps development [113], chatbot development [22], classification of database technology [114], and understanding nonfunctional requirements [115]. To the best of our knowledge, topic analysis of SWSs-related posts on SO is still pending, but SWS is a promising field of software engineering. Therefore, we recognized the need to identify the topics of SWSs, and we utilized BERTopic [93] because of its contextual understanding capacity and high accuracy.
@@ 2.6 Topic Analysis of GitHub Data
Topic modeling with GitHub data involves analyzing text from repositories, issues, pull requests, and comments to identify prevalent themes or topics. This technique can uncover trends, common issues, or popular features within software development communities. Researchers apply algorithms like LDA, BERTopic, and others to categorize the vast amount of unstructured text data on GitHub, helping developers and organizations better understand user needs, project focus, and collaboration patterns. Previous studies have utilized topic modeling to analyze GitHub data, revealing prominent challenges developers face, user requirements, and emerging trends in software projects. This insight assists project managers in task prioritization and resource allocation while promoting streamlined collaboration and well-informed decision-making. By delving into the dynamics of software development communities, researchers contribute to improving software engineering methodologies. For example, Scoccia et al. [27] used LDA topic modeling to identify challenges faced by web developers, correlating real-time GitHub issues with queries on SO to explore common areas in web app development. Similarly, Li et al. [26] applied LDA topic modeling to understand challenges in quantum software engineering by analyzing SO posts and GitHub issue reports. Several other studies [116, 117, 118, 78, 119, 120] have utilized topic modeling with GitHub datasets. Given that GitHub issues and pull requests are often brief and context-specific, BERTopic’s embedding-based approach is particularly well-suited for analyzing this data. Therefore, we chose to utilize BERTopic [93] for our analysis.
@@ 3 Study Design
The main goal of our study is to examine what SWSs developers are asking about. To achieve this goal, we analyzed discussions on SO posts, and issues & pull requests on GitHub. These platforms provide rich datasets, frequently used in research across various domains, including quantum software engineering [26], cryptography APIs [121], concurrency [112], and deep learning [122]. SO offers structured data with questions, answers, and metadata (e.g., accepted answers, scores) but lacks fine-grained topic information related to SWSs. Hence, we first needed to identify posts from SO related to SWSs, group them according to their dominant topic, and then conduct our analysis. For GitHub, we focused on projects related to SWSs identified from our SO data, ensuring consistency across both platforms in the projects we analyzed. Figure 4 shows the overall study design of our research. Below, we described the study design approaches of our work.
@@ 3.1 Stack Overflow Data Collection and Prepossessing
Stack Overflow is a widely used technical Q&A forum among developers. Therefore, we chose it as one of our data sources. Below, we outlined the steps involved in collecting and preprocessing SO data.
Step 1: Gather Scientific Workflow Systems List. Researchers typically use tag-based search functionality on SO to retrieve data, as authors tag their posts with tags (e.g., javascript, python, arrays) to enhance visibility and increase the possibility of receiving answers [123]. However, the field of SWSs spans multiple disciplines (e.g., bioinformatics [124], geospatial analysis [125], software engineering [126], machine learning [127], grid computing [128]), leading to diverse terminologies used by different users. This terminological inconsistency presents a challenge when attempting to conduct searches related to SWSs on SO. Additionally, SWSs have different names (e.g., Galaxy [5], NextFlow [8], SnakeMake [7], Dask [129]) depending on the specific domain or community, further complicating the process of locating these systems through specific tags on SO. Thus, we plan to gather the SWSs list and use it to identify the posts related to them. We gathered a comprehensive list of SWSs by exploring multiple sources, including curated repositories and official documentation [130, 131, 132, 133], to ensure broad coverage across various domains and platforms. Upon merging these lists, we identified overlaps, necessitating the removal of duplicates to ensure the integrity and completeness of our compilation. We also added the web link and GitHub link of each SWS wherever possible. Finally, we listed 352 SWS, which can be accessed using the replication package [53].
@@
We took precautions against potentially noisy data to ensure the accuracy of our dataset and remove any irrelevant information. Despite the possibility of encountering posts with the term celery in their title or body unrelated to the celery workflow system, we conducted a meticulous filtering process. By initially narrowing down our dataset using celery tags, we identified 9,499 posts specifically discussing the celery workflow system. This validation was reinforced by cross-referencing the tags’ descriptions provided by SO. Subsequently, we were left with 139 posts, each potentially containing content unrelated to the celery SWS. These posts varied, with some mentioning the term celery in contexts unrelated to our focus, such as referring to the vegetable, a project name, a part of a filename, or a list ID. To meticulously ensure the precision of our dataset, each of these remaining 139 posts underwent manual scrutiny. Initially, the first author, having over four years of research experience with SWSs and nine years of experience in software development, meticulously reviewed each post, categorizing them as either related or unrelated to the celery workflow system. Following this initial review, the second author, with over eight years of research experience with SWSs, further scrutinized the posts. We only included posts in our final dataset once we all agreed they were relevant. This rigorous process enabled us to identify and exclude six posts unrelated to the celery workflow system (provided in the replication package). Consequently, we finalized our dataset with 9622 posts exclusively pertinent to the celery workflow system.
@@
In some cases, we encountered posts where all content was directly pertinent to SWSs. For instance, in the case of SnakeMake, our extraction process yielded a total of 1,696 posts, meticulously considering tags, Title, and Body. Remarkably, 1,695 of these posts were tagged with snakemake SWS, solidifying their relevance to the SnakeMake SWS. The one remaining post not tagged with SnakeMake also discussed the SnakeMake SWS, further affirming its alignment with our focus. Similar outcomes were observed for KNIME. A thorough investigation led us to identify a grand total of 308 posts related to KNIME, with 305 of them bearing the knime tags. The remaining three, lacking the knime tags, were inherently related to KNIME, underscoring their inclusion in our dataset. Posts that were not explicitly tagged with knime and snakemake underwent the same rigorous manual review process that we applied for celery, ensuring their relevance to the respective SWSs.
@@
Utilizing similar approaches, we explored the Galaxy SWS related posts. Our search encompassed Tags, Title, and Body, yielding a substantial total of 5652 posts. Next, we examined the co-occurring tags to identify potential noise. For the Galaxy dataset, we observed that many posts were also tagged with tags such as android that were unrelated to Galaxy SWS. Consequently, we opted to discard these posts, amounting to the exclusion of 4378 entries. Subsequently, from the remaining 1274 posts, we further refined our dataset by filtering out posts tagged with samsung, mobile, or containing these terms in their title and body. This process led to the removal of 510 additional posts. With 764 posts remaining, we then focused specifically on those related to Galaxy SWS by filtering with the term ansible, associated with Galaxy. This step resulted in the identification of 165 posts relevant to Galaxy. Continuing our scrutiny, we observed 26 posts concerning PHP, 35 posts discussing Java, and 77 posts referencing Galaxy S4-S7. These posts were deemed irrelevant to Galaxy SWS and were consequently excluded. From the remaining 461 posts, following a similar approach discussed for celery, the first author assessed each post, determining their relevance to Galaxy SWS. Subsequently, the second author verified these assessments. Only posts that were unanimously agreed upon as relevant to Galaxy SWS were retained. Ultimately, this rigorous vetting process yielded ten posts directly related to Galaxy SWS. Thus, our finalized dataset comprises a total of 175 posts aligned with Galaxy SWS.
@@
1 Airflow 10602 10589 30168
2 SnakeMake 1696 1696 5945
3 Celery 9628 9622 12512
4 NextFlow 419 418 5075
5 KNIME 308 308 145
6 Galaxy 5648 175 38808
7 Oozie 2066 2052 477
8 Drake 846 846 31186
9 Luigi 362 348 3191
10 Dask 4681 4662 27300
11 NiFi 5187 5187 8311
For airflow, we obtained 10602 posts and we found 10448 posts with airflow SWS tags (This validation was supported by cross-referencing the tag descriptions on SO, confirming that the posts were indeed related to the airflow SWS) and 31 posts having apache in title and body. Then, from the remaining 123 posts we manually checked each posts and found 13 posts unrelated with Airflow SWS. They are mainly talking about heat transfer problems, airflow temperature, respiratory airflow, function name etc. Thus, finally we have 10589 posts related with Airflow SWS.
As there are hundreds of SWS, describing each SWS is beyond our scope, though our approach is identical to that described above. In Table 1, we showed the SWSs, total posts obtained after the query, and total filtered posts. For simplicity, we display the SWSs where we found more than 100 posts for a particular SWS. Our further analysis is also based on the dataset shown in Table 1. These systems are publicly available state-of-the-art systems designed for a broad spectrum of scientific analysis tasks. We gather a total of 41,443 posts. After filtration, we identified 5,540 posts unrelated to the scientific workflow system, and the remaining 35,903 posts are about SWS.
@@ 11 NiFi 5187 5187 8311
Step 3: Preprocessing Scientific Workflow Systems’ posts. We first filtered out the irrelevant information before applying topic modeling techniques. While the post Title concisely summarizes the question, the Body provides crucial context and details that help us accurately identify the topic being discussed. Therefore, in this analysis, we focused on both the post Title and Body. However, the Body content might introduce noise to our analysis, so we removed quotes, HTML tags, links, and any code snippets using regular expressions from the Body of the posts. We also removed stopwords, commonly used words in the English language, such as ’how’, ’can’, and ’at’ that do not significantly affect the meaning of a sentence and can introduce bias. We used the NLTK stopwords list [135] for this operation. Then, we performed lemmatization, which involves reducing a word to its canonical form called a lemma, taking into account the word’s linguistic context (for example, the word ’scale’ is the lemma of the word ’scaling’). All the data pre-processing steps and processed data are provided in the replication package [53]. Below, we provide a concise description of the steps we follow to pre-process the SWS-related SO data.
Removing code snippets and HTML tags: Before applying topic modeling, removing code snippets and HTML tags is essential to ensure cleaner and more meaningful topic extraction. These elements introduce noise by adding structural or syntactic artifacts that do not contribute to the actual discussion. HTML tags, such as <div>, <a href>, and <p>, and programming syntax, like {}, (), ;, :, and if-else statements, can dominate topics if left unprocessed, leading to misleading results. Retaining code snippets may also cause the model to cluster topics based on programming languages or markup structures rather than the underlying challenges being discussed. Removing these elements allowed the topic model to focus on natural language, improving topic coherence and interpretability. We utilized the BeautifulSoup Python library to remove code snippet HTML tags.
def clean_html_and_code(text):
soup = BeautifulSoup(text, "html.parser")
# Remove code blocks
for code_block in soup.find_all("code"):
code_block.decompose()
# Remove HTML tags
cleaned_text = soup.get_text(separator=" ")
return cleaned_text
Stopwords and special character Removal: To improve the quality of our topic modeling process, we applied stopword removal to eliminate commonly used words in the English language that do not significantly contribute to the meaning of a sentence and may introduce bias. For this, we utilized the NLTK stopwords list [135] and further expanded it by appending names of specific SWSs, such as Nextflow, Galaxy, and Snakemake. This refinement ensures that the topic model focuses on the core issues discussed in the dataset rather than being influenced by frequent mentions of specific workflow system names. To implement this process efficiently, we defined a Python function to systematically remove these stopwords from our data, enhancing the clarity and relevance of the extracted topics.
def remove_stopwords(text):
words = text.split()
filtered_words = [word for word in words if word.lower() not in stop_words]
return .join(filtered_words)
[…]
Lemmatization: Lemmatization is the process of converting words to their base form (lemma). One of the most efficient ways to perform lemmatization is by using spaCy [136], a fast and optimized NLP library. To achieve this, we loaded the English language model (en_core_web_sm) and processed the text with spaCy’s pipeline. We transformed each token in the text into its lemma using the .lemma_ attribute. We chose this approach because of its high speed and contextual understanding.
def lemmatize_text(text):
doc = nlp(text)
lemmatized_tokens = [token.lemma_ for token in doc]
return . join(lemmatized_tokens)
3.2 GitHub Data Collection and Preprocessing
To complement our analysis of developer challenges on SO, we also examined issue reports and pull requests from GitHub, a widely used platform for collaborative software development. GitHub repositories provide direct insights into real-world problems encountered during the development and maintenance of SWSs. In this subsection, we described our strategy for collecting and preprocessing GitHub data to enable effective topic modeling and challenge identification.